A train travels at an average speed of 20 mph through urban areas, 50 mph
through suburban areas, and 75 mph through rural areas. If a trip consists of
traveling half an hour through urban areas, 3-1/2 hours through suburban areas,
and 3 hours through rural areas, what is the train’s average speed for the entire
trip?
a. 50 mph
b. 53-2/7 mph
c. 54-3/7 mph
d. 58-4/7 mph
e. 59-2/7 mph
: A train travels at an average speed of 20 mph through urban areas, 50 mph
through suburban areas, and 75 mph through rural areas. If a trip consists of
traveling half an hour through urban areas, 3-1/2 hours through suburban areas,
and 3 hours through rural areas, what is the train’s average speed for the entire
trip?
a. 50 mph
b. 53-2/7 mph
c. 54-3/7 mph
d. 58-4/7 mph
e. 59-2/7 mph

Answer by orca(336)

Remember that the average speed for the entire trip is defined as
average speed=(the total distance traveled)/(the total time spent)
the total distance traveled is:
20*(1/2)+50*(3 1/2)+75*3=10+175+225=410 miles
The total time spent is:
1/2+3 1/2 + 3= 7 hours
So the average speed is:
410/7=58 4/7
Therefore choose option d.